# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v2

@@ A Level Total Order (LTO) on a set $S$ can be defined as follows:
A leveled total order not only provides an ordering across distance groups but also enables detailed, structure-based comparisons within each group.
By defining both inter- and intra-container comparisons, this ordering allows arguments to leverage container-level properties (like distance) and within-container properties (like adjacency) simultaneously. This flexibility is particularly useful in graph theoretic proofs like the SSNC, where it may be necessary to combine distance-based and structural properties.
Defining a leveled total order gives rigor to proofs, as it offers a structured, multi-layered way to approach ordered sets that are partitioned by one property but require further comparability within each partition.
This formalization could also have broader applications beyond the SSNC. It might apply to any problem where elements need to be organized hierarchically but also require nuanced relationships within each level of the hierarchy—such as layered networks, hierarchical clustering, and complex sorting problems.
@@
We previously introduced the concept of containers, which are crucial for understanding the structure of the graph. Interior neighbors are specifically those nodes for which both endpoints are located within the same container. By examining these interior neighbors, we gain insights into how containers relate to one another and how they function internally.
@@
, where $N^{+}(z)$ is the set of out-neighbors of $z$, and $N^{++}(x)$ is the set of second neighbors of $x$. The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
We contrast interior neighbors with exterior neighbors, which reveal how a container interacts with adjacent containers. This interaction is critical for the Seymour Second Neighborhood Conjecture, as the nodes within the current container depend on the connections to the next container. However, it is essential to consider other possibilities that may also influence this relationship.
To better understand the distinction between interior and exterior neighbors, consider Example 5.2. In this example, arcs that connect nodes within the same container (or box) represent interior neighbors. In contrast, arcs that extend to nodes in the next container illustrate exterior neighbors.
@@
Containers are ordered based on their distance from the chosen minimum degree node. Up to this point, all arcs have either been interior or forward-facing. However, it is important to consider the possibility that some arcs may also direct back toward the chosen minimum degree node. Addressing this potential will help us understand how these back arcs could impact the Seymour Second Neighborhood Conjecture.
Now that we have established the definitions of interior and exterior nodes, we can delve into the decreasing sequence property. This property refers specifically to the sizes of the exterior sets within the oriented graph. As the DSP path algorithm progresses, the exterior nodes — which are not part of the neighborhood of the minimum degree node — will systematically decrease in number. This phenomenon reflects a yin and yang relationship between the interior and exterior nodes: while the interior nodes maintain and potentially increase their interconnections, the exterior nodes steadily diminish, ensuring a balanced growth within the graph.
In this dynamic, the minimum degree node serves as a catalyst for this process, driving the algorithm forward by imposing constraints and guiding the restructuring of connections. In contrast, the decreasing size of the exterior nodes acts as a bottleneck, effectively limiting the expansion of the graph’s out-degrees. This interplay is crucial for the success of the DSP path algorithm, as it maintains the structural integrity of the graph while facilitating necessary transformations.
This relationship between the interior and exterior nodes will be the focal point of our proof throughout the rest of this paper, providing a framework for understanding how the DSP path algorithm operates effectively.
With this definition of interior and exterior nodes now introduced, we are now ready to now introduce the
@@ A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
This lemma aligns with the ranked structure we’re building, where each container $C_{i}$ is distinct in terms of reachability and influence. By enforcing this strict order under the absence of back arcs, we’re creating a clear layering within the graph, which further justifies the ”multi-level total order” approach.
Back arcs, if present, could disrupt the unidirectional progression implied by containers, potentially creating feedback loops. By assuming no back arcs, we ensure that each container maintains its place in the total ordering, which adds robustness to the ranked structure we’re developing.
Containers, as defined in the GLOVER framework, serve not only as organizational units but also as bottlenecks that regulate the flow of connectivity within the graph. By grouping nodes based on their shortest-path distance from the minimum degree node $v_{0}$, containers segment the graph into hierarchical levels. The decreasing sequence property ensures that the size of exterior neighborhoods diminishes as we move outward from $v_{0}$, creating a natural bottleneck effect. This reduction in connectivity simplifies the analysis of degree growth and neighbor interactions while preserving the structural integrity of the graph. Such bottlenecks are critical for maintaining the layered structure of the GLOVER framework, particularly in problems like the Seymour Second Neighborhood Conjecture, where second-level connectivity is a key focus.
@@
Let $v_{0}$ be the node with the minimum out-degree in the oriented graph $G$. - By definition, for any node $v\in G$, $d^{+}(v_{0})\leq d^{+}(v)$, where $d^{+}(v)$ is the out-degree of node $v$. - This ordering provides a baseline for comparison because all other nodes are either equal to or have greater out-degree than $v_{0}$.
The analysis begins at $v_{0}$, which serves as the foundation of the GLOVER structure. - All outgoing arcs from $v_{0}$ lead to nodes at the next level (or ”container”) $C_{1}$. - By the properties of the GLOVER structure (e.g., no back arcs), the relationships between nodes are strictly directed, ensuring no circular dependencies.
Assume that for any node $v$ in a container $C_{i}$, the following properties hold: - $d^{+}(v)\geq d^{+}(v_{0})$. - All properties derived from $v_{0}$ apply recursively to nodes in $C_{i}$.
Consider any node $u\in C_{i+1}$ that is connected to a node $v\in C_{i}$. - By the inductive hypothesis, $v$ satisfies all conditions derived from $v_{0}$. - Since $u$ receives its inbound arcs from $v$, and $d^{+}(v)\geq d^{+}(v_{0})$, the relationships in $C_{i+1}$ are consistent with the properties of the Glover structure. - Furthermore, no node $u$ in $C_{i+1}$ can violate the inductive assumptions because $v_{0}$ ensures a well-ordering, propagating constraints downward.
Starting from $v_{0}$, the well-ordering propagates through all containers $C_{i}$, ensuring that: - The out-degree of any node is anchored to $v_{0}$. - There is no upward deviation in degree that contradicts the structure’s constraints.
Thus, the GLOVER structure supports induction by leveraging the minimum degree node $v_{0}$ as an anchor, ensuring consistency throughout the graph.
@@
Lemma 5.4 (Interior Degree Greater than 1) is important because it constrains the out-neighbor structure of the graph when the minimum out-degree $v_{0}$ exhibits the decreasing sequence property. It guarantees that each out-neighbor of $v_{0}$ must have at least one other out-neighbor within $N^{+}(v_{0})$ , which helps to prevent the out-degree of nodes from being too heavily concentrated in the second neighborhood.
The GLOVER structure partitions the neighbors of nodes into interior and exterior categories in a systematic way. This partitioning provides a powerful framework for analyzing oriented graphs, particularly in controlling and understanding degree properties.
A notable consequence of this structured representation is the ability to guarantee that for any oriented graph $G$ represented as a GLOVER, the degree of a chosen minimum out-degree node $v_{0}$ will not double due to cycles in container $C_{1}$. More formally:
Proposition 5.1.
Cycle in Container 1 If $G$ is an oriented graph represented as a GLOVER with $v_{0}$ as the chosen minimum degree node of degree $\delta$, and if there exists a cycle in the first container $C_{1}$ of length $\delta$, then the degree of $v_{0}$ will not double.
@@
This proof directly shows that a contradiction arises if the exterior degree exceeds the out-degree, which would violate the decreasing sequence property. The argument about the doubling of the out-degree in $G^{2}$ is key in demonstrating that the structure defined by the GLOVER representation prevents such a situation, thereby ensuring that $|ext(x,y)|<d^{+}(x)$.
The results presented in 5.1(Cycle in Container 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1, which demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles. While our work is similar to this result, we will extend the framework by leveraging container-based structures to provide a polynomial-time solution, highlighting the interplay between neighboring containers and their shared arcs.
@@
The central concept of this proof revolves around the relationship between decreasing exteriors and back arcs. This relationship is not merely general; it specifically addresses the concern of potential double counting involving second neighbors that arise from back arcs through $y$.
The additional condition - ensuring that a neighbor of $y$ does not also serve as a first neighbor of $x$ is vital. This condition helps to prevent the overlap in counting, particularly in situations where $z$, a second neighbor of $y$, could inadvertently be counted as a first neighbor of $x$.
Given that containers are structured by their distance from the chosen minimum degree node, it follows that if a node $z$ is a second neighbor of $y$ and also a first neighbor of $x$, this necessitates the presence of a back arc from $z$ to a neighbor of $x$. Therefore, the lemma’s validity heavily relies on the configurations of back arcs in our construction, particularly in the context of a potential counterexample.
In summary, the interplay between back arcs and the classification of neighbors is crucial in Lemma 5.6 (Weak Decreasing Exteriors Lemma). By maintaining the distinction between first and second neighbors through careful conditions, we can mitigate the risk of double counting and strengthen the argument presented in the proof.
This next lemma looks into the Weak Decreasing Exteriors Lemma and considers the assumption of back arcs. There is a particular scenario that needs to be evaluated where a second neighbor of $y$ and first neighbor of $z$ can also be a first neighbor of x. The above lemma, called 5.6(Weak Decreasing Exteriors Lemma), has an explicit assumption that gets rid of this. In the next chapter we will do a lot of work to make sure that these back arcs are no longer a problem. However we will consider a second version of this lemma that gets rid of the back arcs assumption.
@@
The theoretical foundation established in this section underpins the subsequent arguments and proofs. By thoroughly examining the properties of oriented graphs, we prepare to tackle the complexities of back arcs and their implications for the overall graph structure.
@@ 6 Back Arcs
In this section, we analyze the role of back arcs in oriented graphs and their effects on graph properties. Understanding back arcs is crucial for assessing how they influence the interior and exterior degrees, as well as the overall structure of the graph.
@@
Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within containers of these counterexamples. Specifically, it shows that all the out-neighbors of interior arcs within a container must have the same degree, which is known as out-degree regularity. If, additionally, the nodes in this container have equal in-degrees, then the container achieves total regularity. It is important to note that this regularity emerges as a counterexample strictly due to the interaction between the Decreasing Sequence Property (DSP) and the minimum degree node $v_{0}$.
In Example 5.5 and Table 1, we examined a minimum degree node $v_{0}$ with degree 6 in a regular graph. The interior degrees of nodes increase outwardly: $v_{0}$ has an interior degree of 0, nodes in $C_{1}$ have 1, nodes in $C_{2}$ have 2, and so forth. This outward growth reflects the structure of the graph, where nodes are grouped into containers based on their distance from $v_{0}$.
A back arc is defined as an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$, effectively pointing against the distance-based order. If such a back arc were introduced, it would imply that the source node $y$ gains new neighbors at a lower level, potentially doubling its degree. This contradicts the hierarchical organization of containers, where each level is strictly defined by shortest-path distances from $v_{0}$.
Challenge:
To better understand why back arcs are not possible in this graph, consider the following questions:
  • Can you construct a back arc in this graph that respects the container construction rules and distance-based ordering from $v_{0}$?
  • If a back arc is introduced, does it lead to contradictions or disrupt the hierarchical structure of containers?
Explanation:
By the construction of the GLOVER framework, all nodes in $C_{i}$ are farther from $v_{0}$ than those in $C_{j}$ for $j<i$. A back arc $(y,z)$ would imply that $z$ is closer to $v_{0}$ than $y$, violating this ordering.
Additionally, introducing a back arc disrupts the unidirectional flow of the graph by creating a feedback loop. This contradicts the fundamental property of containers: that each level represents a discrete, non-overlapping set of nodes ordered by distance from $v_{0}$. Without back arcs, the GLOVER structure preserves a clear layering, enabling rigorous analysis of degree growth and neighbor interactions.
@@
The framework under 6.2 (No Back Arc to Early Container) generalizes the solution to handle all potential back arc scenarios in the graph. Thus, we have strengthened the overall structure, making the algorithm robust against these issues and ensuring that the exterior degree conditions are preserved across all containers.
@@
The concept of a first back arc inherently assumes that no back arcs exist up to a certain point. Lemma 6.2 (No Back Arc to Early Container) supports this assumption, relying on the doubling of nodes’ interior degrees, a condition not necessary for Lemma 6.4. The strength of this lemma lies in its capacity to identify the first back arc; if such an arc cannot be established, it reinforces the graph’s structural integrity. This lemma is founded on our ability to well-order the graph using containers and the Decreasing Sequence Property (DSP). With this ordering, we highlighted regularity within these containers through Lemma 6.1 (Regular Interior Degrees), which proved essential in showing that these back arcs would lead to a node’s degree doubling if a node’s interior degree was already doubling.
@@
The exploration of back arcs reveals their significant impact on the dynamics of oriented graphs. By removing them, we can better navigate the structures of oriented graphs and in particular it paves the way to investigate interior degree dynamics.
@@ 7 Interior Degrees Doubling
This section focuses on the important aspect of interior degree doubling within oriented graphs. We investigate the conditions under which this occurs and the implications for the overall degree distribution of nodes.
@@
The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others. Initially, we accounted for nodes where interior degrees had doubled and ensured consistency when exterior degrees doubled as well. This distinction was critical for our algorithm up to this point. However, with regular containers, interior nodes are now indistinguishable from one another. This means we can swap a node $v_{i}$ with an exterior node $v_{e}$ if their degrees have doubled, ensuring that we have a node whose overall degree has doubled. This flexibility simplifies the process and ensures that the algorithm maintains consistency.
@@
Considering Lemma 7.2 (Interior Degree Double), we observe that it represents a specific instance of the Seymour Second Neighborhood Conjecture. Here, each container functions as its own self-contained oriented subgraph, where we pose the same core question: is there a node whose degree doubles? The affirmative answer we obtain, even when constructing a counterexample, suggests promising progress toward resolving the overall conjecture.
Furthermore, we can expand upon the process illustrated in Lemma 7.2 (Interior Degree Double)). Within the proof, we constructed a monotonically decreasing sequence of subsets, which cannot continue indefinitely. This sequence reflects the essence of the Decreasing Sequence Property (DSP), where we observe that as interior nodes increase, the exterior nodes necessarily decrease. This requirement is what we meant by “decreasing” in DSP: the exterior set gradually diminishes as nodes transition to the interior.
This dynamic can be interpreted as a form of equilibrium, reminiscent of the principle of yin and yang. Each node integrated into the interior necessitates a corresponding reduction in the exterior set, thereby preserving a balance. This interplay ensures that interior growth does not occur in isolation; instead, it aligns directly with constraints on the exterior, maintaining the graph’s structural integrity.
Such interaction is essential for upholding the properties defined by DSP. It underscores the inherent relationship between interior and exterior neighbors, reinforcing the notion that the expansion of one aspect imposes natural limitations on the other.
In summary, examining the doubling of interior degrees within this context provides critical insights into the degree distribution in oriented graphs. This understanding informs the proofs that will be discussed in subsequent sections, enhancing our grasp of both local and global properties of the graph.

### v3

@@ A Level Total Order (LTO) on a set $S$ can be defined as follows:
This order means that we have order in two ways. First, the containers themselves are ordered. Second, the nodes within the containers are ordered. This ordering must also be consistent within the containers.
The notion of a non-triviality requirement is to distinguish this ordering from other traditional total orderings on graphs.
@@
Now that we have introduced containers, we can talk arcs where both nodes are inside the containers, these are interior neighbors. Interior neighbors will be important to our study on Seymour’s Second Neighborhood Conjecture and our search for a counterexample.
In particular we saw things about interior neighbors in 4.3 (Minimum Out-Degree 3 with Neighbors 1). The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1 and we see that when all the neighbors in that subgraph had outdegree of 1, $v_{0}$’s degree did not double.
@@
, where $N^{+}(y)$ is the set of out-neighbors of $y$, and $N^{++}(x)$ is the set of second neighbors of $x$. The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
We contrast interior neighbors with exterior neighbors, which reveal how a container interacts with adjacent containers. This interaction is again critical for the Seymour Second Neighborhood Conjecture, as the nodes within the current container depend on the connections to the next container.
As an exercise to visualize interior and exterior neighbors, consider example 5.2. Exterior arcs are the double arrows between the boxes. Interior arcs are not drawn but can be mentally placed either between any two nodes in the red or blue box.
@@
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. In order to do so though, we need to consider the option of back-arcs and see how they might hurt our algorithm. These have the possibility of altering a container’s distance to the minimum degree node. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture where they may force us to double count nodes.
Having defined our arcs, we can talk about the decreasing sequence property. This property is explaining the nature of how the exterior degrees, and thus the sizes of the containers, decreases as we get further from the minimum degree node. Exterior degrees, though, are a part of a node’s degree equation. So there’s another part of it. As we move further from the minimum degree node, the size of the interior degrees increases. This phenomenon reflects a yin and yang relationship between the interior and exterior degrees.
In all of this, the minimum degree node is serving as a catalyst and driving force for the process. At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s outdegrees.
With this definition of interior and exterior nodes now introduced, we are now ready to now introduce the Graph Level Order (GLOVER) data structure.
@@ A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
What the GLOVER data structure does is transform the partitions and distance metric into an ordering for containers. This ordering is also ordering the nodes within the containers, so the GLOVER data structure is a multi-level total order.
The one thing that could disrupt this total ordering is a back arc. This could potentially lead to a feedback loop. By assuming no back arcs, we ensure that each container maintains its place in the total ordering.
@@
Let $v_{0}$ be the node with the minimum out-degree in the oriented graph $G$. By definition, for any node $v\in G$, $d^{+}(v_{0})\leq d^{+}(v)$, where $d^{+}(v)$ is the out-degree of node $v$. This ordering provides a baseline for comparison because all other nodes are either equal to or have greater out-degree than $v_{0}$.
The analysis begins at $v_{0}$, which serves as the foundation of the GLOVER structure. All outgoing arcs from $v_{0}$ lead to nodes at the next level (or ”container”) $C_{1}$. By the properties of the GLOVER structure (e.g., no back arcs), the relationships between nodes are strictly directed, ensuring no circular dependencies.
Assume that for any node $v$ in a container $C_{i}$, the following properties hold:
  • $d^{+}(v)\geq d^{+}(v_{0})$.
  • All properties derived from $v_{0}$ apply recursively to nodes in $C_{i}$.
Consider any node $u\in C_{i+1}$ that is connected to a node $v\in C_{i}$. By the inductive hypothesis, $v$ satisfies all conditions derived from $v_{0}$. Since $u$ receives its inbound arcs from $v$, and $d^{+}(v)\geq d^{+}(v_{0})$, the relationships in $C_{i+1}$ are consistent with the properties of the Glover structure. Furthermore, no node $u$ in $C_{i+1}$ can violate the inductive assumptions because $v_{0}$ ensures a well-ordering, propagating constraints downward.
Starting from $v_{0}$, the well-ordering propagates through all containers $C_{i}$, ensuring that:
  • The out-degree of any node is anchored to $v_{0}$.
  • There is no upward deviation in degree that contradicts the structure’s constraints.
Thus, the GLOVER structure supports induction by leveraging the minimum degree node $v_{0}$ as an anchor, ensuring consistency throughout the graph. ∎
@@
Lemma 5.4 (Interior Degree Greater than 1) is important because we can see some structural properties in general oriented graphs that may coincide with degree doubling. In particular, we see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree to double.
@@
Lemma 5.5 is very similar to (Exterior Less than Degree), Lemma 5.4 (Interior Degree Greater than 1), with the notable exception that Lemma LABEL:extbd1 refers to any node with the DSP, whereas we only made Lemma 5.4 true for the minimum degree node.
The results presented in 5.4(Interior Degree Greater then 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1. There, he demonstrated that a minimum degree vertex with no cycles in its induced graph has a degree that effectively doubles.
Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an exhaustive search solution with the algorithm we’re developing.
@@
The first thing to notice in this proof are the assumptions. We assume that there are no back arcs and that a node’s interior degree doubles. Both of these assumptions has an important purpose.
First we will talk about the lack of back arcs. We have already talked about how back arcs can be a hindrance to our algorithm overall. This is one of the main points where they can be a hindrance. In particular there exists a way to double count nodes as both neighbors and second neighbors if back arcs exist. We will speak more about this in the next section, but for now we have simply assumed that back arcs do not exist.
Secondly, we assumed that a node’s interior degree doubled. This is similar to the notion of solving the containers separately, then joining them together again to solve the main problem. We are assuming that this not $y$ was in another container and its interior degree doubled in that container. In doing so, $y$ is now in a better position to have its overall degree double, needing only to have its exterior degree double.
@@
We have now defined the data structure that we will use throughout the rest of this paper. In doing so, we have a better understanding of interior and exterior arcs which will help guide us in our search for a counterexample.
@@ 6 Back Arcs
With a theoretical foundation underway, we are ready to launch an investigation into back arcs. We will look into whether these arcs can exist in a counterexample, and what other properties of oriented graphs must be true.
@@
Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within containers of these counterexamples. In particular, it shows that all the out-neighbors of interior arcs within a container must have the same degree.
In Example 5.5 and Table 1, we examined a minimum degree node $v_{0}$ with degree 6 in a regular graph. The interior degrees of nodes increase outwardly: $v_{0}$ has an interior degree of 0, nodes in $C_{1}$ have 1, nodes in $C_{2}$ have 2, and so forth.
A back arc is defined as an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$. If such a back arc were introduced, it would imply that the source node $y$ gains new neighbors at a lower level, potentially doubling its degree. This contradicts the hierarchical organization of containers, where each level is strictly defined by shortest-path distances from $v_{0}$.
@@
Lemma 6.4 (No Back Arc to Early Container) is a special case where we can apply Lemma 6.1 (Regular Interior Degrees). From this, we can see that a back arc to a parent container violates the orientated nature of the graph.
@@
We stated that back arcs could be detrimental to our algorithm, so the removal of them in this section was important. Now we can better navigate the other properties of oriented graphs.
@@ 7 Interior Degrees Doubling
This section introduces the concept of interior degree doubling. This is similar to solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s the interior degree inside of a container. This can be seen as a localized version of Seymour’s Second Neighborhood Conjecture, but for individual containers instead of the overall graph.
@@
The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others. Previously we were searching for a node in a container whose interior degree had doubled. That would be the key. What indistinguishabable tells us is that if one node’s interior degree doubles inside a container, then all the node’s interior degrees double. That means that we can select any of those nodes.
@@
Lemma 7.2 (Interior Degree Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles. By finding a positive answer to this, and not just a positive answer, but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture.
The way the proof was iterated was through both strong induction and a monotonically decreasing sequence of subsets. This cannot continue indefinitely. Once again, we see why the Decreasing Sequence Property (DSP) received its name.
This doubling of interior degrees will help us with the construction of our algorithm in search for a counterexample. It does, however, give us a bit of light into the side of the Conjecture being true.